### v2

# NUTSHELL: A Dataset for Abstract Generation from Scientific Talks

Abstract
Scientific communication is receiving increasing attention in natural language processing, especially to help researches access, summarize, and generate content. One emerging application in this area is Speech-to-Abstract Generation (SAG), which aims to automatically generate abstracts from recorded scientific presentations. SAG enables researchers to efficiently engage with conference talks, but progress has been limited by a lack of large-scale datasets. To address this gap, we introduce NUTSHELL, a novel multimodal dataset of *ACL conference talks paired with their corresponding abstracts. We establish strong baselines for SAG and evaluate the quality of generated abstracts using both automatic metrics and human judgments. Our results highlight the challenges of SAG and demonstrate the benefits of training on NUTSHELL. By releasing NUTSHELL under an open license (CC-BY 4.0), we aim to advance research in SAG and foster the development of improved models and evaluation methods. [^1] https://huggingface.co/datasets/maikezu/nutshell
NUTSHELL: A Dataset for Abstract Generation from Scientific Talks
Maike Züfle1, Sara Papi2, Beatrice Savoldi2, Marco Gaido2, Luisa Bentivogli2, Jan Niehues1 1Karlsruhe Institute of Technology, 2Fondazione Bruno Kessler {maike.zuefle,jan.niehues}@kit.edu, {spapi,bsavoldi,mgaido,bentivo}@fbk.eu
1 Introduction
Abstracts are essential in scientific communication, allowing researchers to quickly grasp the key contributions of a paper. With the ever-growing number of publications, abstracts help researchers stay informed without reading full papers. Beyond their practical utility, abstracts also pose a significant challenge for natural language generation models: abstracts are a specialized form of summarization that not only condenses content but also promotes the work, often using domain-specific terminology and structured language.
Scientific summarization has been widely studied in natural language processing, including summarizing entire articles (Collins et al., 2017; Mao et al., 2022; Liu et al., 2024), particularly in the medical domain (Kedzie et al., 2018; Cohan et al., 2018; Gupta et al., 2021), generating abstracts from citations (Yasunaga et al., 2019; Zanzotto et al., 2020), summarizing specific paper sections (Takeshita et al., 2024), and leveraging knowledge graphs for abstract generation (Koncel-Kedziorski et al., 2019).
Figure 1: NUTSHELL, a dataset for Speech-to-Abstract Generation (SAG) from scientific talks.
With the growing availability of recorded conference talks, a new challenge emerges: generating abstracts from spoken content or Speech-to-Abstract Generation (SAG). The abstracts offer researchers a quick way to assess relevant talks without watching entire recordings. Additionally, as conferences include more virtual content, automatically generated summaries enable efficient engagement with recorded talks (Murray et al., 2010).
While speech summarization has been explored in domains like news (Matsuura et al., 2024), YouTube videos (Sanabria et al., 2018), and meeting minutes (McCowan et al., 2005; Janin et al., 2003), large-scale datasets for scientific talk abstract generation are lacking. Existing work (Lev et al., 2019) aligns transcripts with the corresponding papers and extracts overlapping textual segments as summaries. However, these segments are drawn from the paper rather than the talk itself, failing to capture the distinct contributions, framing, and nuances conveyed in spoken presentations. Other studies have focused on summarizing TED Talks (Koto et al., 2014; Kano et al., 2021; Vico and Niehues, 2022; Shon et al., 2023), which target a broad audience and prioritize inspiration and engagement over technical content.
To bridge this gap, we introduce NUTSHELL a new multimodal dataset for abstract generation from scientific talks. Built from recorded presentations of *ACL conferences, the dataset pairs abstracts with their corresponding spoken content and video, offering a valuable resource for future research. To validate the quality of the abstracts as concise and well-structured summaries of the talks – i.e., capturing the essence of the presentations in a nutshell – we performed a human assessment, which confirmed their effectiveness and suitability for the SAG task.
To establish baselines for SAG using our dataset, we evaluate three model types: (1) a cascaded model combining automatic speech recognition (ASR) with text-based summarization, (2) a state-of-the-art speech-language model (SpeechLLM) without fine-tuning, and (3) a SpeechLLM fine-tuned on our dataset.
Our contributions are three-fold:
  1. 1.
    We introduce NUTSHELL, a novel dataset for abstract generation from scientific talks comprising 1,172 hours, which is released under CC-BY 4.0 License on HuggingFace;1
  2. 2.
    We provide baselines with different model types for comparison in future research, evaluated using both standard automatic metrics (e.g., ROUGE) and the emerging LLM-as-a-judge approach (Shen et al., 2023);
  3. 3.
    We conduct human evaluations to assess the quality of the abstracts and validate the suitability of automatic metrics for the SAG task.
conferences year # examples total audio average audio average words
h min per abstract
train ACL,NAACL, EMNLP 2017-2021 4000 808.3 12.1 $\pm$ 11.2 142.8 $\pm$ 36.1
dev ACL 2022 885 146.4 9.9 $\pm$ 3.6 141.9 $\pm$ 36.5
test EMNLP, NAACL 2022 1431 217.1 9.1 $\pm$ 4.3 147.6 $\pm$ 37.4
total ACL, NAACL, EMNLP 2017-2022 6316 1171.8 11.1 $\pm$ 9.9 143.7 $\pm$ 36.5
Table 1: Dataset statistics for NUTSHELL. The number of words is obtained by splitting the abstract at whitespaces.
2 The NUTSHELL Dataset
In this section, we introduce the new NUTSHELL resource. We chose to build our corpus upon the the ACL Anthology [^2] https://aclanthology.org since it provides a rich collection of multimodal resources (talks and abstracts) and open-access licensing. Starting from 2017, a significant number of papers published in the main *ACL conferences (ACL, EMNLP, and NAACL) include a video of the presentation, all released under the Creative Commons Attribution 4.0 license. This makes *ACL an ideal resource for building a multimodal dataset for the SAG task.
In the following, we present a feasibility assessment of SAG through human evaluation (§2.1). Then, we describe the collection process performed to create NUTSHELL, together with the final dataset statistics (§2.2).
2.1 Are paper abstracts “good” talk summaries?
Before creating the corpus, we establish the validity of our data by investigating whether abstracts represent a good summary of the associated talk. To this aim, we conduct a qualitative check on a data sample of 30 talk-abstract pairs from the ACL Anthology. We involve a total of 5 annotators, who are all domain experts and thus familiar with scientific material. [^3] Annotators include the paper authors and their colleagues. To verify Inter-Annotator Agreement (IAA), a double annotation by different experts was carried out on 15 pairs.
Since we are interested in understanding whether paper abstracts are informative enough to represent a good summary of the talk, we asked evaluators to annotate: (1) Whether the information in the abstract is all uttered by the presenter in the talk; (2) The span of information present in the abstract that was not contained in the talk, if any; (3) Whether the abstract summarizes all important information presented in the talk. The human evaluation procedure, including the annotation template, is described in App. A.
The results indicate that $70.0\%$ of the abstracts are considered good summaries by annotators as they contain important information about the talk. However, $63.3\%$ of the abstracts also contain information not explicitly present in the talk itself. To better understand this, we conducted a qualitative analysis of the annotated spans corresponding to this missing information. We found that these spans typically involved dataset names, model names, shared task references (e.g., evaluation campaigns), or URLs (e.g., link to the resource or model being released). Notably, these elements are often displayed on slides but not explicitly verbalized by presenters. [^4] This issue could be overcome by exploiting the videos, as this information is typically shown in the slides. While out of scope for SAG, NUTSHELL includes the videos, making it a useful resource also for more complex multimodal tasks.
Despite this issue, the evaluation of automatic models against the same ground truth abstract can be considered fair, as models are equally penalized by this category of missing information. Moreover, it is worth noting that establishing a single ground truth for summarization tasks is still an open challenge (Zhang et al., 2024), given the inherent variability in human-produced summaries.
Both, questions (1) and (3) have an inter-annotator agreement of $\kappa=0.466$, indicating moderate agreement (Landis and Koch, 1977), which can be regarded as acceptable given the subjective nature of evaluating summaries. While criterion (3) naturally involves subjective judgments about information importance, the lower agreement on criterion (1) can also be attributed to borderline cases, where small phrasing differences were sometimes overlooked by individual annotators. Such subtleties led to occasional discrepancies in annotator decisions, but were manually reviewed.
In summary, the manual evaluation confirmed both the feasibility of the SAG tasks and, despite the noted challenges, the overall reliability and usefulness of our resource.
2.2 Collection and Dataset Statistics
We collected talks from 16 ACL Anthology events: 6 ACL, 6 EMNLP, and 4 NAACL, including workshops, shared tasks and industry tracks. For each paper (both long and short format), we extracted the video and the associated abstract already available on the paper website. We exclude papers with invalid URLs, videos without audio, or abstracts missing from the paper page. Additional details on the data collection can be found in App. B.
Lastly, we split the dataset into training (years 2017 to 2021), dev (ACL 2022), and test (EMNLP/NAACL 2022). These splits reflect a realistic evaluation setup, where models are trained on past data and tested on the most recent, unseen examples. In total, the corpus contains 1,172 hours of audio content corresponding to 6,316 different presentations. Full statistics are reported in Table 1.
3 Analysis
To demonstrate the quality and usability of our corpus, as well as provide baselines for future works, we develop and evaluate four different models using both automatic metrics and human evaluation.
3.1 Experimental Setting
3.1.1 Models
To establish baselines for the SAG task, we analyze the performance of four models described as follows. Prompts, model, generation, and additional training details are provided in App. C.
Whisper + LLama3.1-8B-Instruct.
A cascaded solution, where the audio is first transcribed with openai/whisper-large-v3 (Radford et al., 2022), and then meta-llama/Llama-3.1-8B-Instruct (Dubey et al., 2024) is prompted to generate the abstract from the generated transcript.
Model RougeL BERTScore Llama3.1-7B-Instruct Human (on subset)
F1 $\uparrow$ F1 $\uparrow$ Score with Expl. $\uparrow$ Plain Score $\uparrow$ Avg. Rank $\downarrow$ Avg. Rank $\downarrow$
Whisper + LLama3.1-8B-Instruct 22.14 86.62 77.84 82.47 1.24 1.53
Qwen2-Audio-7B-Instruct 15.02 84.65 45.57 36.81 3.43 2.87
End2End Finetuned 23.89 86.66 68.78 73.53 1.98 1.60
End2End Zero-Shot 16.08 84.13 45.97 39.90 3.35 N/A
Table 2: We report results on the NUTSHELL test set for four models: a cascaded approach (Whisper+Llama-3.1-8B-Instruct), an existing SpeechLLM (Qwen2-Audio), and an end-to-end HuBERT+QFormer+Llama3.1-8B-Instruct model, either finetuned on our data (End2End Finetuned ) or trained on audio instruction-following data (End2End Zero-Shot). Avg. Rank, assigned by an LLM judge or human annotators, reflects the mean ranking per model.
Qwen2-Audio-7B-Instruct.
The Qwen/Qwen2-Audio-7B-Instruct (Chu et al., 2024) model, an existing SpeechLLM [^5] By SpeechLLM, we refer to the combination of a speech encoder and an LLM through a learned modality adapter (Gaido et al., 2024)., which is used out of the box without any fine-tuning.
End2End Zero-Shot.
A SpeechLLM composed of HuBERT (Hsu et al., 2021) as speech encoder, meta-llama/Llama-3.1-8B-Instruct as LLM, and a QFormer (Li et al., 2023) as adapter. The SpeechLMM is built to handle long audio inputs (App. C) and obtained by training only the adapter in two steps: (a) contrastive pretraining (Züfle and Niehues, 2024) to align the LLM representations for the speech and text modalities using MuST-C (Di Gangi et al., 2019) and Gigaspeech (Chen et al., 2021), and (b) fine-tuning on instruction-following tasks, including ASR, speech translation, and spoken question answering using MuST-C and Spoken-SQuAD (Lee et al., 2018). Therefore, the model is not trained or fine-tuned on NUTSHELL and operates in zero-shot for the SAG task.
End2End Finetuned.
A SpeechLLM trained using the same contrastive pretraining procedure as End2End Zero-Shot but subsequently fine-tuned on our NUTSHELL dataset. This not only evaluates the direct impact of task-specific datasets on the SAG performance, but it also ensures the feasibility of the task and the suitability of the collected data.
3.1.2 Evaluation
Metrics.
We use standard (text) summarization metrics: ROUGE (Lin, 2004) – a text similarity metric that has been widely adopted for LM evaluation (Grusky, 2023) that focuses on n-gram overlap between the hypothesis and reference –, and BERTScore (Zhang et al., 2020) – a neural-based metric that measures the pairwise similarity of contextualized token embeddings between the summary and its reference. Also, we rely on LLM-as-a-judge (Shen et al., 2023; Zheng et al., 2024) where the LLM [^6] We use Llama-3.1-8B-Instruct (Dubey et al., 2024) as the judge using the prompts reported in Fig. 2 in Section D.2. is prompted to assign a score to each output, using the reference abstract as context (Score with Expl.). The score is based on four criteria: (1) relevance, (2) coherence, (3) conciseness, and (4) factual accuracy.777(1) Does the predicted abstract capture the main points of the gold abstract?, (2) Is the predicted abstract logically organized and easy to follow?, (3) Is the predicted abstract free from unnecessary details?, (4) Are the claims in the predicted abstract consistent with the gold abstract? We also report results where the LLM judge provides a single score without explanations (Plain Score), as well as results where it ranks the given abstracts instead of scoring them individually (Avg. Rank).
All these metrics have known limitations and no metric is conclusively best for evaluating the SAG task: both ROUGE and BERTScore are known to fail to fully capture the extent to which two summaries share information (Deutsch and Roth, 2021) while LLM-as-a-judge is sensitive to prompt complexity and the length of input (Thakur et al., 2024) and struggle to distinguish similar candidates (Shen et al., 2023). For this reason, we complement the automatic scores with human evaluation.
Human Evaluation.
For the human evaluation, nine annotators – all experts in the field – were provided with the generated abstracts and the ground truth abstract. We use the same randomly sampled 30 test set examples as in Section 2.1 and validate their representativeness, which is discussed in App. E. Each sample is evaluated by three annotators. They follow the same criteria as the LLM evaluation but rank models instead of assigning scores. Detailed instructions are in App. E. As the End2End Zero-Shot model performance was comparable to that of Qwen2-Audio – also being a zero-shot model – and given that Qwen2-Audio is an established SpeechLLM with a distinct architecture, we exclude the End2End Zero-Shot from this analysis.
3.2 Results
Automatic Evaluation.
Table 2 presents the performance of our models on the NUTSHELL test set. Among them, the cascaded model (Whisper + Llama3.1-8B-Instruct) achieves the highest scores across all LLM-based evaluation metrics. Instead, looking at both n-gram- and neural-based metrics, the End2End Finetuned model achieves the highest RougeL and BERTScore. In addition, Qwen2-Audio and our End2End Zero-Shot models demonstrate similar performance across all automatic metrics, showing a noticeable gap compared to the other two models. These results highlight the importance of our dataset for building high-performing end-to-end models, as the substantial gap between the cascaded and End2End Zero-Shot models is effectively bridged through fine-tuning on the NUTSHELL dataset.
For a more granular analysis, Table 3 in Section D.2 provides results for the LLM-based metrics. Given that all models except Qwen2-Audio rely on Llama3.1-8B-Instruct, one might question whether the Llama-based judge could introduce bias in favor of these models. To address this, we perform additional evaluations using Qwen/Qwen2-7B (Yang et al., 2024) as the judge (Table 4 in Section D.2), which confirm the same ranking, eliminating any concerns about evaluator bias.
Human Evaluation.
As shown in Table 2, the human evaluation results closely align with the LLM-based judgments: the cascaded model ranks first, followed closely by the finetuned model while Qwen2-Audio ranks last. Notably, the gap between the first two models is small, whereas the difference between the second and third models is substantial – consistent with the LLM-based evaluation. This suggests that automatic metrics reliably capture both subtle and large performance differences between models. IAA, measured using pairwise rankings (Bojar et al., 2016) reached $\kappa=0.53$, which is acceptable given the close ranking of the top two systems.
4 Conclusion
In this work, we introduce NUTSHELL, a novel dataset for SAG from recorded *ACL conference talks. By releasing this dataset under an open license, we hope to foster further advancements in SAG research and encourage the development of more effective models and evaluation techniques. Future work could explore the integration of the video content provided in the corpus, offering an additional modality for enriching the generation process and further improving abstract quality.
5 Limitations
While the current study provides a new resource and offers valuable insights about the SAG task, two main limitations should be noted:
  • The analysis focused on the speech-to-text abstract generation task. However, our dataset also provides access to the corresponding videos, which were not utilized here. Future research could explore the integration of video content as an additional modality to enhance the generation process and improve the quality of the abstracts.
  • The human evaluation was limited in scope, involving only a small set of models and samples. Future work could expand this evaluation to include more models and a larger number of samples to better assess the performance of different metrics and determine which is most effective in various contexts.
Potential Risks
Generating automatic summaries for scientific talks carries the risk that automatic summaries may misrepresent key findings or lack scientific accuracy. However, we hope that by providing more high-quality training data, summarization models can be improved and lead to more reliable and accurate summaries.
Acknowledgments
This work has received funding from the European Union’s Horizon research and innovation programme under grant agreement No 101135798, project Meetween (My Personal AI Mediator for Virtual MEETtings BetWEEN People). We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid ACC Cyfronet AGH for providing computer facilities. Beatrice Savoldi and Marco Gaido are supported by the PNRR project FAIR - Future AI Research (PE00000013), under the NRRP MUR program funded by the NextGenerationEU.
We thank Leonard Bärmann, Béni Egressy, Mia Fuß, Lukas Hilgert, and Danni Liu for their contributions to the data annotation process.
Appendix A Human Evaluation: Are abstracts good summaries of the talk?
Figure 3: Instructions for annotators to evaluate whether the paper abstracts are good and informative abstracts for the ACL talks.
We aim to assess whether paper abstracts can serve as effective abstracts for *ACL talks. To this end, we conducted a human evaluation by randomly sampling 30 examples from our dataset. The annotation team consisted of five individuals (four women and one man), including the paper authors and their colleagues. All annotators were already familiar with the NLP domain, scientific presentation and writing, and the task itself. They are experts in Natural Language Processing, holding at least a master’s degree in NLP or a related field, with some holding PhDs or professorial positions. Their ages ranged from 25 to 55.
The annotation guidelines were initially developed by the authors and subsequently refined in collaboration with the annotators to ensure a shared and well-defined set of evaluation criteria. Detailed instructions for the human annotators are provided in Fig. 3. The annotation template also included a comment section for uncertain cases, though no comments were submitted.
Appendix B Dataset Details
We include all *ACL conferences from 2017 to 2022 in NUTSHELL, covering main conferences, Findings, industry tracks, and workshops. As not all conferences are held every year, the number of talks varies accordingly. Table 5 provides a detailed overview.
Split Conference Year Talks
Train ACL 2017 140
2018 185
2019 244
2021 849
EMNLP 2017 93
2018 221
2021 1480
NAACL 2018 120
2019 114
2021 554
Dev ACL 2022 885
Test EMNLP 2022 465
NAACL 2022 966
Table 5: Number of talks per conferences in the NUTSHELL dataset.
Appendix C Baseline Details
Generation Settings
We evaluate four different models to establish baselines for abstract generation from spoken ACL talks. The evaluations were conducted on a single NVIDIA A100-SXM4-40GB GPU.
For all models, we use the default generation parameters and apply greedy search, following the usage instructions for meta-llama/Llama-3.1-8B-Instruct [^8] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct (Dubey et al., 2024), Qwen/Qwen2-Audio-7B-Instruct [^9] https://github.com/QwenLM/Qwen2-Audio (Chu et al., 2024) and the contrastively pretrained models from Züfle and Niehues (2024) [^10] https://github.com/MaikeZuefle/contr-pretraining.
Cascaded Model
For the cascaded model, we segment the audio into 30-second chunks and transcribe them using openai/whisper-large-v3 (Radford et al., 2022). The transcribed chunks are then concatenated and processed by meta-llama/Llama-3.1-8B-Instruct (Dubey et al., 2024) to generate the abstract. Inference took 5:40 hours on a single NVIDIA A100-SXM4-40GB GPU, including transcribing and summarizing.
Since the model’s outputs often included a title and category for the talk, we explicitly prompt it to generate only the abstract. This adjustment was not necessary for the other models.
We use the following prompt:
System Prompt:
[…]
Prompt:
[…]
Qwen2-Audio
For Qwen/Qwen2-Audio-7B-Instruct (Chu et al., 2024), inference took 50 minutes on a single NVIDIA A100-SXM4-40GB GPU. We use the system prompt as provided in the code documentation9.
System Prompt:
You are a helpful assistant.
Prompt:
[…]
Model RougeL BERTScore Llama3.1-7B-Instruct
F1 $\uparrow$ F1 $\uparrow$ Score with Expl. $\uparrow$ Plain Score $\uparrow$ Avg. Rank $\downarrow$
Whisper + LLama31-Instruct 23.26 86.81 77.75 84.30 1.23
Qwen2-Audio 16.26 84.94 48.42 39.50 3.47
End2End Finetuned 24.47 86.71 70.67 75.73 1.83
Table 6: Baseline Results, the finetuned model is a HuBERT + Qformer + LLama31Instruct model on the subset used for human annotation (30 examples).
Contrastively Pretained Models
For the contrastively pretrained model, we follow Züfle and Niehues (2024) and adopt their settings10, including training configurations, hyperparameters, and system prompts. The SpeechLLM consists of HuBERT (Hsu et al., 2021) as speech encoder, meta-llama/Llama-3.1-8B-Instruct as LLM, and a QFormer (Li et al., 2023) as adapter. We choose HuBERT as an encoder in contrast to the bigger and more powerful openai/whisper-large-v3 (Radford et al., 2022), as it needs less memory and is therefore more suitable for the summarization task of longer audio. However, due to the extended duration of the audio inputs, we additionally introduce two modifications:
  1. 1.
    We segment the audio into one-minute chunks, encode each chunk using the encoder and then concatenate the encoded representations before passing them through the adapter and LLM backbone.
  2. 2.
    We use a batch size of 1 for fine-tuning with NUTSHELL.
Despite these adjustments, we encountered memory limitations for audio files exceeding 35 minutes. In such cases, we truncate the audio to 35 minutes, which affects one example in the test set.
The training of the models was conducted on four NVIDIA A100-SXM4-40GB GPUs. The contrastive pretraining took 33 hours on four GPUS. Finetuning on ASR, speech translation, and spoken question answering data took 30 hours, finetuning on the NUTSHELL dataset took 2:10 hours. Generating the outputs of the test set (on a single NVIDIA A100-SXM4-40GB GPU) took 2:35 hours.
System Prompt:
[…]
Prompt:
Summarize this talk to create an abstract for an ACL Paper:
Appendix D Evaluation Details
We evaluate the results of our models using automatic metrics including ROUGE, BERTScore, and LLM-as-a-judge.
D.1 ROUGE and BERT Score
As automatic metrics, we use ROUGE111111 (Lin, 2004) and BERTScore (Zhang et al., 2020). Concretely, we compute ROUGE-L, which focuses on the longest common subsequence, with DD/sacrerouge (Deutsch and Roth, 2020), as recommended by Grusky (2023) and for BERTScore, we use the bertscore implementation from HuggingFace [^12] https://huggingface.co/spaces/evaluate-metric/bertscore and report the F1-score.
D.2 LLM as a judge
To evaluate the model outputs, we also use an LLM as a judge, specifically meta-llama/Llama-3.1-8B-Instruct (Dubey et al., 2024). The LLM assigns a score to each output using the reference abstract as context, based on four criteria: (1) relevance (Does the predicted abstract capture the main points of the gold abstract?), (2) coherence (Is the predicted abstract logically organized and easy to follow?), (3) conciseness (Is the predicted abstract free from unnecessary details?), and (4) factual accuracy (Are the claims in the predicted abstract consistent with the gold abstract?). Additionally, we report results where the LLM provides a single overall score without explanations and results where it ranks the given abstracts instead of scoring them individually. The prompts are given in Fig. 2. If the model fails to return a valid json dictionary, we instead take the first number after the score name in the output. We present the results for all four criteria, the average score, the score without explanations, and the ranking in Table 3. One potential concern is that this LLM might be biased, as all our models except Qwen2-Audio are based on Llama-3.1. However, we find this is not the case. When using Qwen/Qwen2-7B (Yang et al., 2024) as the judge, we obtain the same ranking as with Llama. The results with Qwen-as-a-judge can be found in Table 4.
Appendix E Human Evaluation for Model Outputs
We evaluate the models using ROUGE (Lin, 2004), BERTScore (Zhang et al., 2020), and LLM-as-a-judge. However, it is known that automatic evaluation metrics can come with limitations. Namely, the first two metrics may not fully capture semantic overlap (Deutsch and Roth, 2021), while LLM-as-a-judge is sensitive to prompt phrasing (Thakur et al., 2024) and struggles to distinguish between closely similar candidates (Shen et al., 2023). To validate the reliability of our automatic evaluation scores and better understand model behavior, we complement these metrics with a human evaluation. This allows us also to verify the robustness of our findings.
Specifically, we asked nine domain experts (four women and five men) to rank model outputs relative to the reference abstract, with each example annotated by three independent annotators. All annotators were already familiar with the NLP domain, scientific writing and presentation, and the task itself. They are experts in Natural Language Processing, holding at least a master’s degree in NLP or a related field, with some holding PhDs or professorial positions. Their ages ranged from 25 to 55. The annotation instructions are provided in Fig. 4.
We conduct this human evaluation on a randomly selected subset of 30 test examples. We consider this subset representative, as the model rankings based on automatic metrics remain consistent with those on the full test set. The corresponding automatic scores for this subset are reported in Table 6. We want to include three diverse models in our human evaluation: a zero-shot model, a cascaded model, and a model finetuned on our dataset. Since we have two zero-shot models (Qwen2-Audio and our contrastively pretrained zero-shot model) that perform similarly, we decided to exclude one for efficiency in the human evaluation. We keep the Qwen2-Audio model as this is an already existing and widely used SpeechLLM.
Figure 4: Instructions for human annotators for ranking model outputs.